Given that right angled  triangle  PQR angle Q=90deg
Angle p=25.3DEG AND QR=7.4cm
Calculate  the length of PR

Grade:10

1 Answers

Sk
3 years ago
In ∆PQR,
sin P = QR/PR
sin 25.3°   = 7.4/PR
0.42=7.4/PR
PR =7.4/0.42=17.61 Answer
Thus, the length of hypotenuse or PR can be determined by using trigonometric function sine for  angle P .